OpenStudy (anonymous):

What is the function that describes this sequence? x y 1 3 2 6 3 9 4 12 a)y = 3/n b)y = 3n c)y = n/3 d)y = n + 3

OpenStudy (snowsurf):

Its b. Because everything has a common multiple of 3.

OpenStudy (jdoe0001):

\(\begin{array}{llll} x&y \\\hline\\ 1&3&3\cdot x\to 3\cdot 1\\ 2&6&3\cdot x\to 3\cdot 2\\ 3&9&3\cdot x\to 3\cdot 3\\ 4&12&3\cdot x\to 3\cdot 4 \end{array}\)